# Computing almost commuting bases of ODOs and Gelfand-Dikey hierarchies

### v1

@@ Abstract
Almost commuting operators were introduced in 1985 by George Wilson to present generalizations the Korteweg-de Vries hierarchy, nowadays known as Gelfand-Dikey (GD) hierarchies. In this paper, we review the formal construction of the vector space of almost commuting operators with a given ordinary differential operator (ODO), with the ultimate goal of obtaining a basis by computational routines, using the language of differential polynomials. We use Wilson’s results on weigheted ODOs to guarantee the solvability of the triangular system that allows to compute the homogeneous almost commuting operator of a given order in the ring of ODOs. As a consequence the computation of the equations of the GD hierarchies is obtained without using pseudo-differential operators. The algorithms to calculate the almost commuting basis and the GD hierarchies in the ring of ODOs are implemented in SageMath, and explicit examples are provided.
MSC[2020]: 13N10, 13P15, 12H05
@@ 1 Introduction
Almost commuting operators were introduced in 1985 by George Wilson in [20] to present the Lax representations and generalizations of the Korteweg-de Vries hierarchy, nowadays known as the Gelfand-Dikey (GD) hierarchies [5, 6]. The generalization of the Korteweg-de Vries hierarchy, passing from a second-order linear differential operator to an $n$-th order operator $L$ was suggested by Krichever, Gelfand and Dickey, see references in [5, 6]. The standard construction of these generalized hierarchies is based on fractional powers $L^{m/n}$ of an operator $L$ of order $n$, which live in a ring of pseudo-differential operators.
The question addressed is this paper is the design of a symbolic algorithm that allows the computation of the GD hierarchies but does not require computations with pseudo-differential operators. We develop an algorithm that requires computations in a ring of differential operators $\mathcal{D}$ whose coefficients belong to a ring of differential polynomials. This is possible by the notion of almost commuting operator defined by G. Wilson in  [21], as an operator $A\in\mathcal{D}$ such that $[L,A]=LA-LA$ has order $\leq n-2$. We reorganized previously known ideas (by Wilson and many others) to prove that the vector space $W(L)$ of almost commuting operators with a given ordinary differential operator $L$
$W(L)=\{A\in\mathcal{D}\mid\rm ord[L,A]\leq n-2\}=Z((L))_{+}.$
equals the set $Z((L))_{+}$ of positive parts of the elements of the centralizer $Z((L))$ of $L$ in the ring of pseudo-differential operators, see Theorem 3.5.
Our motivation comes from the need of having a basis of almost commuting operators, and explicit formulas for integrable hierarchies, to determine differential operators with nontrivial centralizers. Hence our ultimate goal is to obtain such bases by computational routines. We saw in the work of Grunbaum [8] the possibility of developing symbolic algorithms. We used preliminary implementations of these algorithms in Maple to compute the examples in previous works of some of the authors, see [13, 15], and [16]. We need meaningfulexamples of prime order ODOs to extend the results in [16], and then develop a suitable Picard-Vessiot theory for spectral problems.
@@
Wilson defined a weight function on the ring of differential polynomials =$\mathcal{R}=\mathbb{Q}\{U\}$, that guaranteed the existence of a unique basis $\mathcal{H}(L)=\{P_{m}\mid m\in\mathbb{N}\}$ whose operators are homogeneous with respect to the weight function. A natural approach to seek for an homogeneous almost commuting operator $P_{m}$ of order $m$ is to define the formal operator
$\tilde{P}_{m}=\partial^{m}+y_{2}\partial^{m-2}+\ldots+y_{m-1}\partial+y_{m}$
whose coefficients are differential variables in $Y=\{y_{2},\ldots,y_{m}\}$ with respect to $\partial$, and force $[L,\tilde{P}_{m}]$ to be a differential operator of order $\leq n-2$. This produces a differential system $\mathbb{E}_{m,n}$ of $m-2$ equations given by differential polynomials in $\mathbb{Q}\{U,Y\}$. We proved that $\mathbb{E}_{m,n}$ is a triangular system w.r.t. the variables $Y$ in Theorem 4.5 and moreover that
@@
are differential polynomials, homogeneous with respect to an extended weight function on $\mathbb{Q}\{U,Y\}$. We use Wilson’s results on weigheted ODOs to guarantee the solvability of this triangular system, that allows to compute a unique solution set $Z$ of $\mathbb{E}_{m,n}$ w.r.t. $Y$. We can solve $\mathbb{E}_{m,n}$ by using the integration methods for differential polynomials described in  [1] and  [2], combined with backwards substitution.
@@
The algorithms to compute the almost commuting basis and the GD hierarchies, in the ring of ODOs are implemented in the SageMath package dalgebra [9, 19]. More details of this implementation and of the computations performed are included in Section 6.
The paper is organised as follows. In Section 2 we provide all the formal definitions for the differential algebra concepts that are used throughout the paper. We recall in Section 3 the main structural result of Wilson for almost commuting operators. In Section 4 we will show the direct method for computing almost commuting bases. First we will present how to obtain a triangular differential system for solving this differential system. We proceed to show in Section 5 how this method produces the first elements on the GD hierarchies for an operator $L$ of order 3 and 5. Finally, Section 6 contains a description of the software implemented in SageMath [19], and the results obtained with our symbolic algorithms.
@@ 2 Preliminaries
A differential ring $(R,\partial)$ is a ring $R$ (assumed associative, with identity) equipped with a specified derivation $\partial$, that is, an additive map $\partial:R\rightarrow R$ satisfying the product rule: $\partial(xy)=\partial(x)y+x\partial(y)$ for all $x,y\in R$. The elements $r\in R$ such that $\partial(r)=0$ are called constants of $R$, and the collection of these constants forms a subring of $R$. A differential field $(K,\partial)$ is a differential ring which is a field.
@@
The notation $y^{\prime}$, $y^{\prime\prime}$, $y^{\prime\prime\prime}$ will be also used for $y^{(i)}$, $i=1,2,3$. Observe that $R\{y\}$ is also a differential ring with the natural extended derivation, that we also call $\partial$ abusing notation. We can iterate this construction, adding several differential variables. Given differential variables $y$ and $z$ over $R$, the ring of differential polynomials in $y$ and $z$ is denoted by $R\{y,z\}$ and defined as $R\{y\}\{z\}\simeq R\{z\}\{y\}$.
A differential ring $(R,\partial)$, defines a ring of (linear) differential operators $R[\partial]$, which is a non commutative ring, where the multiplication of two elements of is the composition of the two differential operators, defined by the rule $\partial r=\partial(r)+r\partial$ for all $r\in R$. It is an Ore polynomial ring [7]. A differential operator $A\in R[\partial]$ can be written in a unique way as
$A=a_{n}\partial^{n}+a_{n-1}\partial^{n-1}+\ldots+a_{1}\partial+a_{0},\mbox{ with }a_{i}\in R,\,\,\,\mbox{ with }a_{n}\neq 0.$
Then the order of $A$ is $n$, denoted $\rm ord(A)=n$, with the convention $\rm ord(0)=-\infty$. This order has the same property as the degree of a polynomial: the order of a product is the sum of the orders and the order of a sum is bounded by the maximal order of the summands. The coefficient of the highest order term of $A$ is the leading coefficient of $A$, denoted $\operatorname{lc}(A)=a_{n}$. We say that an operator $A\in R[\partial]$ of order $n$ is in normal form if it is monic, i.e., it has $\operatorname{lc}(A)=1$, and the coefficient of the order $n-1$ term is zero, namely,
@@
In [12, 22] these operators are called elliptic operators. Differential operators with coefficients in a differential field can be always brought to normal form, by using a change of variable and conjugation by a function, which are the only two possible automorphisms.
Let us assume that the ring of constants of $R$ is a field $\mathbf{C}$ of zero characteristic. Since $R[\partial]$ is a non-commutative ring, we can define the corresponding Lie-bracket over $R[\partial]$ as a $\mathbf{C}$-bilinear map [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] defined by
$[A,B]=AB-BA.$
It is clear that $A$ and $B$ commutes if and only if $[A,B]=0$. We also say that $[A,B]$ is the commutator of $A$ and $B$. We define the centralizer of $A$ in $R[\partial]$ as
$\mathcal{Z}(A)=\{B\in R[\partial]\ :\ [A,B]=0\}.$
Since [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is a $\mathbf{C}$-bilinear map, then $\mathcal{Z}(A)$ is a $\mathbf{C}$-module.
@@
Let $A,B$ be two differential operators in $R[\partial]$ of orders $n$ and $m$ respectively. Then $\rm ord([A,B])\leq n+m-1$. If $A$ and $B$ are both in normal form, then $\rm ord([A,B])\leq n+m-3$.
@@
Since $\operatorname{lc}(AB)=\operatorname{lc}(A)\operatorname{lc}(B)=\operatorname{lc}(BA)$, we have that the term of order $n+m$ of $[A,B]$ is always zero, implying $\rm ord([A,B])\leq n+m-1$. Moreover, if $A$ and $B$ are in normal form, we can write $A=\partial^{n}+\tilde{A}$ and $B=\partial^{m}+\tilde{B}$ where $\tilde{A}$ and $\tilde{B}$ have orders bounded by $n-2$ and $m-2$ respectively. Then
$[A,B]=[\partial^{n},\partial^{m}]+[\tilde{A},\partial^{m}]+[\partial^{n},\tilde{B}]+[\tilde{A},\tilde{B}].$
The first bracket always vanishes and the orders of the other three brackets are bounded by $n+m-3$. ∎
Let us denote by $R((\partial^{-1}))$ the ring of pseudo-differential operators in $\partial$ with coefficients in the differential ring $R$, defined as in [7]
$R((\partial^{-1}))=\left\{\sum_{i=-\infty}^{n}a_{i}\partial^{i}\mid a_{i}\in R,n\in\mathbb{Z}\right\},$
where $\partial^{-1}$ is the inverse of $\partial$ in $R((\partial^{-1}))$, $\partial^{-1}\partial=\partial\partial^{-1}=1$, and it satisfies the following commutation rule with elements $r\in R$
@@
Given $A=\sum_{i=-\infty}^{n}a_{i}\partial^{i}$, with $a_{n}\neq 0$ then $n$ is called the order of $A$, denoted by $\rm ord(A)$ and $a_{n}$ is called the leading coefficient of $A$. We define the positive part of $A$ as the differential operator
@@
We denote by $A_{-}=A-A_{+}$. Given $A,B\in R((\partial^{-1}))$ of orders $m$ and $n$ respectively, we can also define their commutator $[A,B]$. Using the natural decomposition $A=A_{+}+A_{-}$, we can see that
$[A,B]=[A_{+},B_{+}]+[A_{-},B_{+}]+[A_{+},B_{-}]+[A_{-},B_{-}],$
@@ 3 Almost Commuting Operators
Given a monic $n$-th order differential operator $L$ in $R[\partial]$, it is well known that it has a unique monic $n$-th root, denoted by $L^{1/n}$, see for instance [7, Proposition 2.7]. In addition, $L^{1/n}$ determines the centralizer in $R((\partial^{-1}))$ of $L$, denoted by $\mathcal{Z}((L))$ and equal to
$\mathcal{Z}((L))=\left\{\sum_{j=-\infty}^{m}c_{j}(L^{1/n})^{j}\mid c_{j}\in{\bf C},m\in\mathbb{Z}\right\},$ (3)
by [7, Theorem 3.1]. This implies that $\mathcal{Z}((L))$ is commutative. This result is a generalization of a famous theorem by I. Schur in [17], originally stated for analytic coefficients, to differential operators with coefficients in a differential ring $R$. This theorem has a long history, see for instance [7, Sections 3 and 4], and the references therein.
@@
Let us consider a monic differential operator $L\in R[\partial]$ of order $n$. For every $A\in\mathcal{Z}((L))_{+}$ it holds that $\rm ord([L,A])\leq n-2$.
@@
Given $B\in\mathcal{Z}((L))$. Let us consider the natural decomposition $B=B_{+}+B_{-}$ and observe that $B_{-}$ is a pseudo differential operator of order at most $-1$. Since $B\in\mathcal{Z}((L))$ then $0=[L,B]=[L,B_{+}]+[L,B_{-}]$ implying that the order of $[L,B_{+}]$ equals the order of $[L,B_{-}]$. This order is less than or equal to $n-2$. ∎
@@
Let $L\in R[\partial]$ be a differential operator of order $n$. We say that $A\in R[\partial]$ almost commutes with $L$ if $\rm ord([L,A])\leq n-2$. We denote by $W(L)$ the set of all differential operators that almost commute with $L$.
@@
$[L,cB_{1}+B_{2}]=c[L,B_{1}]+[L,B_{2}].$
Thus, it is clear, by the properties of the order, that $cB_{1}+B_{2}\in W(L)$ again, proving the first statement.
Given $A\in W(L)$, let us prove that $\operatorname{lc}(A)$ is a constant. If $\rm ord(A)=m$, let $a_{m}=\operatorname{lc}(A)$ and it is easily proved that the leading term of $[L,A]$ is $na_{m}^{\prime}\partial^{n+m-1}$. Since $\rm ord([L,A])\leq n-2$ then $a_{m}^{\prime}=0$, proving that $a_{m}\in\mathbf{C}$. ∎
Lemma 3.1 implies that $\mathcal{Z}((L))_{+}\subseteq W(L)$. We will prove next that equality holds. Given an $n$-th root $L^{1/n}$ of $L$ we denote its $m$-th power by $L^{m/n}$.
@@
Let us consider a monic differential operator $L\in R[\partial]$ of order $n$ and choose an $n$-th root $L^{1/n}$ of $L$. Then the set $\mathcal{B}(L)=\{B_{j}=(L^{j/n})_{+}\mid j\in\mathbb{N}\}$ is a basis of $\mathcal{Z}((L))_{+}$ as a $\mathbf{C}$-vector space.
@@
The existence of $B=L^{1/n}$ is guarantied by [7], Proposition 2.7. Note that $B$ has order one. Thus $B_{0}=1$ and $\rm ord(B_{j})=j$, for every $\0j\in\mathbb{N}\backslash\{0\}$. Thus $\mathcal{B}(L)$ is a set of linearly independent differential operators and, by (4) and (3), it is a generating set of $\mathcal{Z}((L))_{+}$ as a $\mathbf{C}$-vector space. ∎
@@
Assuming that $L$ has order $n$, we consider the $\mathbf{C}$-basis $\mathcal{B}(L)=\{B_{j}=(L^{j/n})_{+}\mid j\in\mathbb{N}\}$ of $\mathcal{Z}((L))_{+}$ as in Proposition 3.4. By Lemma 3.3, given $A\in W(L)$ of order $m$ then there exists $c_{m}\in\mathbf{C}$ such that $A-c_{m}B_{m}$ belongs to $W(L)$ and has order less than or equal to $m-1$. Now by induction on the order of $A$, we conclude that $A\in\mathcal{Z}((L))_{+}$. ∎

### v2

# Computing almost commuting bases of ODOs and Gelfand-Dickey hierarchies

@@ Abstract
Almost commuting operators were introduced in 1985 by George Wilson to present generalizations of the Korteweg-de Vries hierarchy, nowadays known as Gelfand-Dickey (GD) hierarchies. In this paper, we review the formal construction of the vector space of almost commuting operators with a given ordinary differential operator (ODO), with the ultimate goal of obtaining a basis by computational routines, using the language of differential polynomials. We use Wilson’s results on weighted ODOs to guarantee the solvability of the triangular system that allows to compute the homogeneous almost commuting operator of a given order in the ring of ODOs. As a consequence, the computation of the equations of the GD hierarchies is achieved without using pseudo-differential operators. A new package in SageMath called dalgebra has been designed to perform symbolic calculations in differential domains. The algorithms to calculate the almost commuting basis and the GD hierarchies in the ring of ODOs are implemented in SageMath, and explicit examples are provided.
pacs:
MSC[2020]]13N10, 13P15, 12H05
@@ 1 Introduction
Almost commuting operators were introduced in 1985 by George Wilson in [1] to present the Lax representations and generalizations of the Korteweg-de Vries hierarchy, nowadays known as the Gelfand-Dickey (GD) hierarchies [2, 3]. The generalization of the Korteweg-de Vries hierarchy, passing from a second-order linear differential operator to an $n$-th order operator $L$ was suggested by Krichever, and implemented by Gelfand and Dickey, see references in [2, 3]. The standard construction of these generalized hierarchies is based on fractional powers $L^{m/n}$ of a differential operator $L$ of order $n$, which live in a ring of pseudo-differential operators.
The question addressed in this paper is the design of a symbolic algorithm that allows the computation of the GD hierarchies but does not require computations with pseudo-differential operators. We develop an algorithm that requires computations in a ring of differential operators $\mathcal{D}$, whose coefficients belong to a ring of differential polynomials. This is made possible by the notion of almost commuting operator defined by G. Wilson in  [1], as an operator $A\in\mathcal{D}$ such that [L,A]=LAAL[L,A]=LA-{AL} has order $\leq n-2$. We reorganized previously known ideas (by Wilson and many others) to prove that the vector space $W(L)$, of almost commuting operators with a given ordinary differential operator $L$
$W(L)=\{A\in\mathcal{D}\mid\operatorname{ord}[L,A]\leq n-2\}={Z((L))_{+},}$
equals the set $Z((L))_{+}$, of positive parts of the elements of the centralizer $Z((L))$ of $L$ in the ring of pseudo-differential operators, see Theorem 3.5.
Our motivation comes from the need of having a basis of almost commuting operators, and explicit formulas for integrable hierarchies, to search for new families of algebro-geometric differential operators, those having centralizers that are not polynomial rings and containing pairs of ODOs of coprime orders [4, 5, 6]. Recursive methods exist to compute almost commuting bases, for specific values of the order $n$ of $L$. For $n=2$, recursive formulas for the Korteweg-de Vries hierarchy were given in [7], that can be obtained by means of a recursion operator [8]. For $n=3$ recursive formulas can be found in [9], related with Boussinesq systems. Our ultimate goal is to compute almost commuting bases by computational routines for an arbitrary value of $n$. We saw in the work of Grünbaum [10] the possibility of developing symbolic algorithms for this purpose. We used preliminary implementations of these algorithms in Maple to compute examples of algebro-geometric differential operators in previous works of some of the authors: see [11, 12], and [6]. New examples of prime order ODOs will allow one to extend the results in [6], and then develop a suitable Picard-Vessiot theory for spectral problems. In addition, an operator $A$ of the almost commuting basis of $L$ provides a Lax representation $L_{t}=[A,L]$ and is a preparatory step towards the computation of recursion operators as in [13].
@@
Wilson defined a weight function on the ring of differential polynomials $\mathcal{R}=\mathbb{Q}\{U\}$, that guaranteed the existence of a unique basis $\mathcal{H}(L)=\{P_{m}\mid m\in\mathbb{N}\}$ of $W(L)$, whose operators are homogeneous with respect to the weight function. A natural approach to seek for an homogeneous almost commuting operator $P_{m}$ of order $m$ is to define the formal operator
$\tilde{P}_{m}=\partial^{m}+y_{2}\partial^{m-2}+\ldots+y_{m-1}\partial+y_{m},$
whose coefficients are differential variables in $Y=\{y_{2},\ldots,y_{m}\}$ with respect to $\partial$, and force $L,\tilde{P}_{m}]$ to be a differential operator of order $\leq n-2$. This produces a differential system $\mathbb{E}_{m,n}$ of m2m-2 equations given by differential polynomials in $\mathbb{Q}\{U,Y\}$. We proved that $\mathbb{E}_{m,n}$ is a triangular system w.r.t. the variables $Y$ in Theorem 4.6 and moreover that
@@
are differential polynomials, homogeneous with respect to an extended weight function on $\mathbb{Q}\{U,Y\}$. We use Wilson’s results on weighted ODOs to guarantee the solvability of this triangular system, that allows to compute a unique solution set $Z$ of $\mathbb{E}_{m,n}$ w.r.t. $Y$. We can solve $\mathbb{E}_{m,n}$ by using the integration methods for differential polynomials described in  [14] and  [15], combined with backwards substitution.
@@
A new package in SageMath called dalgebra has been designed to perform symbolic calculations in differential domains. The algorithms to compute the almost commuting basis and the GD hierarchies in the ring of ODOs, are implemented in the the module almost_commuting.py  [16, 17]. More details on this implementation and the computations performed are included in Section 6. A data set has been generated including almost commuting bases and GD hierarchies for orders $n=2,3,5,7$ of $L$, considering it would be useful to researchers in Mathematics and Physics.
The paper is organised as follows. In Section 2 we provide definitions for the differential algebra concepts that are used throughout the paper. We review in Section 3 the main structural result of Wilson for almost commuting operators. In Section 4 we define the system $\mathbb{E}_{m,n}$ that allows the computation of almost commuting bases. We will prove that this system is a triangular differential system with a unique weighted solution $Z$. We proceed to show in Section 5 how the GD hierarchies are obtained as a by-product, and compute them for an operator $L$ of orders $3$ and $5$. Finally, Section 6 contains a description of the implementation in SageMath [17] of the symbolic algorithms to compute almost commuting bases and GD hierarchies.
@@ 2 Preliminaries
A differential ring $(R,\partial)$ is a ring $R$ (assumed associative, with identity) equipped with a specified derivation $\partial$, that is, an additive map $\partial:R\rightarrow R$ satisfying the product rule: $\partial(xy)=\partial(x)y+x\partial(y)$ for all $x,y\in R$. The elements $r\in R$ such that $\partial(r)=0$ are called constants of $R$, and the collection of these constants forms a subring of $R$. A differential field $(K,\partial)$ is a differential ring that is a field.
@@
The notation $^{\prime}, y′′y^{\prime\prime}$, $y^{\prime\prime\prime}$ will also be used for $y^{(i)}$, $i=1,2,3$. Observe that $R\{y\}$ is a differential ring with the natural extended derivation, that we also call $\partial$ abusing notation. We can iterate this construction, adding several differential variables. Given differential variables $y$ and $z$ over $R$, the ring of differential polynomials in $y$ and $z$ is denoted by $R\{y,z\}$ and defined as $R\{y\}\{z\}\simeq R\{z\}\{y\}$.
A differential ring $(R,\partial)$, defines a ring of (linear) differential operators $R[\partial]$, which is a non-commutative ring, an Ore polynomial ring [20], where the multiplication of two elements is the composition of the two differential operators, that is defined by the rule $\partial r=\partial(r)+r\partial$ for all $r\in R$. A differential operator $A\in R[\partial]$ can be written in a unique way as
$A=a_{n}\partial^{n}+a_{n-1}\partial^{n-1}+\ldots+a_{1}\partial+a_{0},\mbox{ with }a_{i}\in R,\,\,\,a_{n}\neq 0.$
Then the order of $A$ is $n$, denoted $\operatorname{ord}(A)=n$, with the convention $\operatorname{ord}(0)=$-$\infty$. This order function shares properties with the degree of a polynomial: the order of a product is the sum of the orders and the order of a sum is bounded by the maximal order of the summands. The coefficient of the highest order term of $A$ is the leading coefficient of $A$, denoted $\operatorname{lc}(A)=a_{n}$. We say that an operator $A\in R[\partial]$ of order $n$ is in normal form if it is monic, i.e. it has $\operatorname{lc}(A)=1$ and the coefficient of the term of order n1n-1 is zero, namely,
@@
In [21, 22] these operators are called elliptic. Differential operators with coefficients in a differential field can be always brought to normal form, by using a change of variable and conjugation by a function, which are the only two possible automorphisms.
Let us assume that the ring of constants of $R$ is a field $\mathbf{C}$ of zero characteristic. Since $R[\partial]$ is a non-commutative ring, we can define the corresponding Lie-bracket on $R[\partial]$ as a $\mathbf{C}$-bilinear map $\cdot$$\cdot$] defined by
[A,B]=ABBA.[A,B]=AB-BA.
It is clear that $A$ and $B$ commute if and only if $[A,B]=0$. We also say that $[A,B]$ is the commutator of $A$ and $B$. We define the centralizer of $A$ in $R[\partial]$ as
$\mathcal{Z}(A)=\{B\in R[\partial]\ \mid[A,B]=0\}.$
Since $\cdot$$\cdot$] is a $\mathbf{C}$-bilinear map, then $\mathcal{Z}(A)$ is a $\mathbf{C}$-module.
The next lemma states a basic property of ordinary differential operators, that for operators in normal form will be important to develop the main results of this paper, for this reason a proof is included.
@@
Let $A,B$ be two differential operators in $R[\partial]$ of orders $n$ and $m$ respectively. Then $\operatorname{ord}([A,B])\leq n+m-1$. If $A$ and $B$ are both in normal form, then $\operatorname{ord}([A,B])\leq n+m-3$.
@@
Since $\operatorname{lc}(AB)=\operatorname{lc}(A)\operatorname{lc}(B)=\operatorname{lc}(BA)$, the term of order $n+m$ of $[A,B]$ is always zero, implying $\operatorname{ord}([A,B])\leq n+m-1$. Moreover, if $A$ and $B$ are in normal form, we can write $A=\partial^{n}+\tilde{A}$ and $B=\partial^{m}+\tilde{B}$ where $\tilde{A}$ and $\tilde{B}$ have orders bounded by n2n-2 and m2m-2 respectively. Then
$A,B]=[\partial^{n},\partial^{m}]+[\tilde{A},\partial^{m}]+[\partial^{n},\tilde{B}]+[\tilde{A},\tilde{B}$].
The first bracket always vanishes and the orders of the other three brackets are bounded by n+m3n+m-3. ∎
Let us denote by $R((\partial^{-1}))$ the ring of pseudo-differential operators in $\partial$ with coefficients in the differential ring $R$, defined as in [20] ( see also [23], Chapter 5 and [2])
$R((\partial^{-1}))=\left\{{\sum_{-\infty<i\leq n}}a_{i}\partial^{i}\mid a_{i}\in R,n\in\mathbb{Z}\right\},$
where $\partial^{-1}$ is the inverse of $\partial$ in $R((\partial^{-1}))$$\partial^{-1}\partial=\partial\partial^{-1}=1$, and it satisfies the following commutation rule with elements $r\in R$
@@
Given $A\in R((\partial^{-1}))$, with $a_{n}\neq 0$ then $n$ is called the order of $A$, denoted by $\operatorname{ord}(A)$ and $a_{n}$ is called the leading coefficient of $A$. The positive part of $A$ is defined as the differential operator
@@
As it is ususal, denote by $A_{-}=A-A_{+}$. Given $A,B\in R((\partial^{-1}))$ of orders $m$ and $n$ respectively, their commutator $[A,B]$ is also defined and using the decomposition $A=A_{+}+A_{-}$, we can see that
$A,B]=[A_{+},B_{+}]+[A_{-},B_{+}]+[A_{+},B_{-}]+[A_{-},B_{-}],$
@@ 3 Almost Commuting Operators
Given a monic $n$-th order differential operator $L$ in $R[\partial]$, it is well known that it has a unique monic $n$-th root, which is the first order pseudo-differential operator with leading term $\partial$, denoted by $L^{1/n}$, see for instance  [2, Proposition 1.3.8] and  [20, Proposition 2.7]. In addition, $L^{1/n}$ determines the centralizer in $R((\partial^{-1}))$ of $L$, denoted by $\mathcal{Z}((L))$ and equal to
$\mathcal{Z}((L))=\left\{\sum_{-\infty<j\leq n}c_{j}(L^{1/n})^{j}\mid c_{j}\in{\bf C},j\in\mathbb{Z}\right\},$ (3)
by [20, Theorem 3.1]. This implies that $\mathcal{Z}((L))$ is commutative. This result is a generalization of a famous theorem by I. Schur in [24], originally stated for analytic coefficients, to differential operators with coefficients in a differential ring $R$. This theorem has a long history, see for instance [20, Sections 3 and 4], and the references therein.
@@
Let us consider a monic differential operator $L\in R[\partial]$ of order $n$. For every $A\in\mathcal{Z}((L))_{+}$ it holds that $\operatorname{ord}([L,A])\leq n-2$.
@@
Given $A\in\mathcal{Z}((L))_{+}$ then $A=B_{+}$, with $B\in\mathcal{Z}((L))$. Let us consider the natural decomposition $B=B_{+}+B_{-}$ and observe that $B_{-}$ is a pseudo-differential operator of order at most 1-1. Since $B\in\mathcal{Z}((L))$ then $0=[L,B]=[L,B_{+}]+[L,B_{-}]$ implying that the order of $L,B_{+}]$ equals the order of $L,B_{-}]$. This order is less than or equal to n2n-2. ∎
@@
Let $L\in R[\partial]$ be a differential operator of order $n$. We say that $A\in R[\partial]$ almost commutes with $L$ if $\operatorname{ord}([L,A])\leq n-2$. We denote by $W(L)$ the set of all differential operators that almost commute with $L$.
@@
$L,cB_{1}+B_{2}]=c[L,B_{1}]+[L,B_{2}].$
Thus, by the properties of the order, $cB_{1}+B_{2}\in W(L)$, proving the first statement.
Given $A\in W(L)$, let us prove that $\operatorname{lc}(A)$ is a constant. If $\operatorname{ord}(A)=m$, let $a_{m}=\operatorname{lc}(A)$. It is easily proved that the leading term of $[L,A]$ is $na_{m}^{\prime}\partial^{n+m-1}$. Since $\operatorname{ord}([L,A])\leq n-2$ then $a_{m}^{\prime}=0$, proving that $a_{m}\in\mathbf{C}$. ∎
Lemma 3.1 implies that $\mathcal{Z}((L))_{+}\subseteq W(L)$. We will next prove that equality holds. Given the monic $n$-th root $L^{1/n}$ of $L$ we denote its $m$-th power by $L^{m/n}$.
@@
Let us consider a monic differential operator $L\in R[\partial]$ of order $n$ and its monic $n$-th root $L^{1/n}$. Then the set $\mathcal{B}(L)=\{B_{j}=(L^{j/n})_{+}\mid j\in\mathbb{N}\}$ is a basis of $\mathcal{Z}((L))_{+}$ as a $\mathbf{C}$-vector space.
@@
[…]
@@
Assuming that $L$ has order $n$, we consider the $\mathbf{C}$-basis $\mathcal{B}(L)=\{B_{j}=(L^{j/n})_{+}\mid j\in\mathbb{N}\}$ of $\mathcal{Z}((L))_{+}$ as in Proposition 3.4. By Lemma 3.3, given $A\in W(L)$ of order $m$ then there exists $c_{m}\in\mathbf{C}$ such that $A-c_{m}B_{m}$ belongs to $W(L)$ and has order less than or equal to m1m-1. Now by induction on the order of $A$, we conclude that $A\in\mathcal{Z}((L))_{+}$. ∎